# Identification and estimation of mediational effects of longitudinal modified treatment policies

### v1

@@ 4.3 Results
Table 2: Most important effect modifiers for the direct and indirect effect of invasive mechanical ventilation on survival through acute kidney failure, as measured by the slope of a simple linear regression of the efficient influence function. Variables with negative slopes indicate populations with lower effects (calculated as intervention minus baseline; positive values indicate that delaying intubation is relatively less harmful/more beneficial).
First, we note that this analysis is exploratory; some subsets may have relatively small sample sizes and we have not calculated standard errors. However, interstitial lung disease (ILD) is the most important variable for effect modification for both the indirect effect through AKI and the direct effect through other mechanisms. The negative effect modification parameter for ILD in Table 2 implies that delaying IMV in patients with ILD has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients without ILD, whereas the positive effect modification parameter in Table 2 means that delaying IMV in patients with ILD has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without ILD. The former effect can be explained by the fact that delaying intubation may be riskier in patients with ILD; the latter effect (smaller in size) might be due to the fact that patients with ILD are more likely to experience mortality due to lung failure rather than AKI.
Likewise, Table 2 shows that delaying IMV in patients with cancer, chronic obstructive pulmonary disease (COPD), high lymphocyte count, costovertebral angle tenderness (CVA), coronary artery disease (CAD), high creatinine values, and those who used home oxygen, has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without those conditions. Delaying IMV in patients who have cirrhosis or who are current smokers has a worse effect on survival (i.e., is more harmful/less beneficial) due to AKI than in patients without those conditions.
The results in Table 2 show that delaying intubation in patients who are immunosuppressed, use home oxygen, or are white, has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients outside those categories. Lastly, delaying intubation in patients with high creatinine values, asthma, CVA, cirrhosis, high bilirubin, and who are former smokers, has a better effect on survival (i.e., is less harmful/more beneficial) due to mechanisms other than AKI than in patients without those conditions.
4.4 Mediation and treatment decisions
By contrast, one could attempt to estimate the total effect for all patients and exclude any patients with a negative value, but this shuts out the possibility of additional decision-making input from physicians, and it ignores the possibility of remediation of the indirect effects (i.e., that additional interventions could be put in place to mitigate the probability of AKI for patients under intubation).
5 Discussion
Supplementary Materials for
S1 Technical note on notation
S2 Identification (Theorem 1)
Proof By the law of total expectation, we can write
$\theta=\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))]$
$=\mathsf{E}[\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))\mid J(\bar{A}^{\mathbbm{d}^{\star}})]]$
$=\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))\mid\bar{J}(\bar{A}^{\mathbbm{d}^{\star}})=\bar{m}]*\mathsf{P}(\bar{J}(\bar{A}^{\mathbbm{d}^{\star}})=\bar{m})$
$=\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{m})\mid\bar{J}(\bar{A}^{\mathbbm{d}^{\star}})=\bar{m}]*\mathsf{P}(\bar{J}(\bar{A}^{\mathbbm{d}^{\star}})=\bar{m})$
$=\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{m})]*\mathsf{P}(\bar{M}(\bar{A}^{\mathbbm{d}^{\star}})=\bar{m})$
Given policies $\mathbbm{d}^{\prime}$ and $\mathbbm{d}^{\star}$, we let $\varphi(\bar{m})=\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{m})]$ and $\lambda(\bar{m})=\mathsf{P}(\bar{M}(\bar{A}^{\mathbbm{d}^{\star}})=\bar{m})$ and will show that $\varphi$ and $\lambda$ are equivalent to the expressions given in the theorem.
First, for $\varphi$, we consider the relabeled dataset where for all $t\leq\tau$, $(L_{t},Z_{t})=(C_{2t-1},C_{2t})$ and $(A_{t},M_{t})=(E_{2t-1},E_{2t})$. In this case, the expression $Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{m})$ is the expected value of the counterfactual outcome of the longitudinal modified treatment policy setting the values of $E$ to the corresponding values $\bar{A}^{\mathbbm{d}^{\prime}}$ and $\bar{m}$. Similarly to Díaz et al. (2023b), this expression is identified by the following strategy.
Let $\mu_{2\tau+1}=Y$ and for $t=2\tau,\dots,1$ recursively define
where $H_{t}$ denotes all variables prior to $E_{t}$. Then by integration, $\mathsf{E}[Y(\mathsf{E}^{\mathbbm{d}^{\prime}})]=\mathsf{E}[\mu_{1}(E_{1}^{\prime},C_{1})]$.
Noting that the desired intervention on $\bar{m}$ is not stochastic (and thus the mediation component of the intervention $\mathbbm{d}^{\prime}$ can be passed into the $\mathsf{Q}$-functions as the future histories $\underaccent{\bar}{m}_{t}$); thus one can see that substituting variables is sufficient to recognize, when $\mathbbm{d}^{\prime}$ is the intervention setting $E$ to the values of $\bar{A}^{\mathbbm{d}^{\prime}}$ and $\bar{m}$, the relation $\mathsf{E}[\mu_{1}(E_{1}^{\prime},C_{1})]=\mathsf{Q}_{L,0}^{\prime}(\bar{m})$, as in the Theorem.
For $\lambda(\bar{m})=\mathsf{P}(\bar{M}(\bar{A}^{\mathbbm{d}^{\star}})=\bar{m})$, we again relabel the data, this time grouping $L,Z,M$ as $S_{t}=(L_{t},Z_{t-1},\bar{M}_{t-1})$ (and again suitably altering the meaning of $H_{t}$ to denote all variables before $A_{t}$, or $H_{A,t}$ in the original labeling). Then $\bar{M}(\bar{A}^{\mathbbm{d}\star})$ is simply a projection of the outcome $S_{\tau+1}$ under the longitudinal modified treatment policy $\bar{A}^{\mathbbm{d}\star}$. Denoting this projection as $S_{t}^{m}=\bar{M}_{t-1}$, we can write $\mathsf{P}(\bar{M}(\bar{A}^{\mathbbm{d}\star})=\bar{m})=\mathsf{E}[\mathbbm{1}(S_{\tau+1}^{m}=\bar{m})]$ and we can again apply a similar identification strategy.
As a shorthand, let $\mathbbm{1}_{\tau+1}^{\bar{m}}=\mathbbm{1}(S_{\tau+1}^{m}=\bar{m})$. Let $\rho_{\tau+1}=\mathbbm{1}_{\tau+1}^{\bar{m}}$ and for $t=\tau,\dots,1$ recursively define
First, by definition we have
$\rho_{\tau}(a_{\tau},h_{\tau})=\mathsf{E}[\mathbbm{1}_{\tau+1}^{\bar{m}}\mid A_{\tau}=a_{\tau},H_{t}=h_{\tau}]$
$=\mathbbm{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{P}(M_{\tau}=m_{\tau}\mid A_{\tau}=a_{\tau},H_{t}=h_{\tau})$
$=\mathbbm{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{E}[\mathbbm{1}(M_{\tau}=m_{\tau})\mid A_{\tau}=a_{\tau},H_{t}=h_{\tau}]$
By definition,
$=\mathbbm{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{Q}_{M,\tau}^{*}(\underaccent{\bar}{m}_{\tau})$
Then for $t=\tau-1$, we similarly have
$=\mathsf{E}[\mathbbm{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{Q}_{M,\tau}^{*}(\underaccent{\bar}{m}_{\tau})\mid A_{\tau-1}=a_{\tau-1},H_{t}=h_{\tau-1}]$
$=\mathbbm{1}(S_{\tau-1}^{m}=\bar{m}_{\tau-2})\mathsf{E}[\mathbbm{1}(M_{\tau-1}=m_{\tau-1})\mathsf{Q}_{M,\tau}^{*}(\underaccent{\bar}{m}_{\tau})\mid A_{\tau-1}=a_{\tau-1},H_{t}=h_{\tau-1}]$
$=\mathbbm{1}(S_{\tau-1}^{m}=\bar{m}_{\tau-2})\mathsf{Q}_{M,\tau-1}^{\star}(\underaccent{\bar}{m}_{\tau-1})$
Repeated iterations yield the pattern
$\rho_{t}(a_{t},h_{t})=\mathbbm{1}(S_{t}^{m}=\bar{m}_{t-1})\mathsf{Q}_{M,t}^{\star}(\underaccent{\bar}{m}_{t})$
S3 IPW identification
Díaz et al. (2021) gives, for a general longitudinal modified treatment policy $\bar{A}^{\mathbbm{d}}$ and outcome $Y$ (without explicit mediation or intermediate confounders):
$\mathsf{E}[Y(\bar{A}^{\mathbbm{d}})]=\mathsf{E}\left[Y\prod_{t=1}^{\tau}\frac{g_{t}^{\mathbbm{d}}(a_{t}|h_{A,t})}{g_{t}(a_{t}|h_{A,t})}\right]$

### v3

@@ 6 Acknowledgments
S8 Simulation study
To test the accuracy of our estimator, we undertook a simulation study described by the following data-generating distribution.
$L_{1}$ $\sim\mathcal{N}(0,1)$
$A_{1}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5L_{1}))$
$Z_{1}$ $\sim\mathcal{N}(0.5(-L_{1}+A_{1}-0.5),1)$
$M_{1}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5(L_{1}-A_{1}-Z_{1}+0.5)))$
$L_{2}$ $\sim\mathcal{N}(0.5(-L_{1}+A_{1}+Z_{1}-M_{1}),1)$
$A_{2}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5(L_{1}-A_{1}-Z_{1}+M_{1}+L_{2})))$
$Z_{2}$ $\sim\mathcal{N}(0.5(-L_{1}+A_{1}+Z_{1}-M_{1}-L_{2}+A_{2}-0.5),1)$
$M_{2}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5(L_{1}-A_{1}-Z_{1}+M_{1}+L_{2}-A_{2}-Z_{2}+0.5)))$
$Y$ $\sim\mathcal{N}(0.5(-L_{1}-A_{1}-Z_{1}-M_{1}+L_{2}+UA_{2}-Z_{2}+VM_{2}),1)$
where $U,V\in{-1,1}$ are parameters that vary across simulation scenarios. We vary the sample size $n\in(1000,5000)$. Each combination of $U,V,n$ is run for 300 replicates.
We consider two treatment policies; $\mathsf{d}$ is the status quo (according to the data-generating process), and $\mathsf{d}^{\prime}$ is the intervention that sets $A_{1}=1$ and then samples $A_{2}$ naturally, according to the data-generating process. Our estimand of interest is $\theta=\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},J(\bar{A}^{d})]$ (see main paper for an explanation of this notation). The true value of the target parameter is $\theta(U,V)=(U-1.25+V/2)/2$.
Estimation was performed using the lcmmtp package with cross-fitting ($K=3$ folds) and an ensemble of learners in the Super Learner framework (van der Laan et al., 2007) including mean, GLM, multivariate adaptive regression splines, and random forest. For each parameter combination $(n,U,V)$, 300 simulation replicates were conducted.
$n$ $U$ $V$ $\theta$ $n\cdot\text{MSE}$ Coverage
$1\cdot 10^{3}$1 -1 -1.375 3.872 0.950
$5\cdot 10^{3}$1 -1 -1.375 4.039 0.930
$1\cdot 10^{3}$1 1 -0.875 3.255 0.963
$5\cdot 10^{3}$1 1 -0.875 3.797 0.953
$1\cdot 10^{3}$ 1 -1 -0.375 4.486 0.920
$5\cdot 10^{3}$ 1 -1 -0.375 3.827 0.943
$1\cdot 10^{3}$ 1 1 0.125 3.296 0.947
$5\cdot 10^{3}$ 1 1 0.125 3.591 0.963
Table S3: Summary of simulation scenarios and results. The table presents the sample size ($n$), parameters $U$ and $V$, the true parameter value ($\theta$), the product of the sample size and mean squared error ($n\cdot\text{MSE}$), and the coverage probability for each scenario in the simulation study.
[…]
Figure S3: Coverage Probability for the parameter estimate across different sample sizes and scenarios. Lines connect points with the same $U$ and $V$ values.
As shown in the figures and table, the method demonstrates good accuracy (with steady $n\cdot\text{MSE}$ across sample sizes) and approximately nominal coverage. As a final remark, we note that we have chosen a relatively simple data-generating process, making use of linear functions, in order to facilitate the calculation of the theoretical value of $\theta$, but those functional forms are not assumed by the model or estimators, which can accommodate much more complex data-generating processes.
Code for the simulation study can be found at https://github.com/bjg345/lcmmtp_sim/.